OpenStudy (anonymous):

If y varies inversely as x and y = 5 when x = 2, find x when y is 4.

OpenStudy (lgbasallote):

\[\large y = \frac{k}{x}\] \[5 = \frac{k}{2}\] \[k = 10\] can you do the rest?

OpenStudy (anonymous):

x=5

OpenStudy (lgbasallote):

you got it

OpenStudy (anonymous):

Thanks :D

OpenStudy (lgbasallote):

oops wait no... it's \[4 = \frac{10}{x}\] \[x = \frac{10}{4}\] sorry

OpenStudy (anonymous):

so x = 2.5???

OpenStudy (lgbasallote):

2.5 or 5/2